OpenStudy (anonymous):

A population of bacteria is modeled by the exponential function P(t) = 45 x 2^(t/3) where t is in days. To the nearest day, when will the population reach 400?

OpenStudy (anonymous):

it's means that P(t)=400 then: 400=45 * 2^(t/3) 2^(t/3)=400/45=80/9 (t/3) ln2 = 80/9 t=80/(3ln2) use calculator to get your answer :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

Hey can you see if you can help me in this too...

OpenStudy (anonymous):

give me 10 min - I need to finish other problem first